Having trouble with systems of equations using elimination. y=2x+3 and y=-x-4
Found 2 solutions by richwmiller, ozybloke:
Answer by richwmiller(9135)

They are both equal to y.
Set them equal to each other.
2x+3=-x-4
3x=-7
x=-7/3
y=7/3-4
y=7/3-12/3
y=-5/3
or we could
y=2x+3 and
y=-x-4
subtract the second from the first
0=3x+7
-7=3x
-7/3=x
That looks familiar.

Answer by ozybloke(7)

y = 2x + 3
y = -x - 4
Use the substitution method
If y = 2x + 3 then replace the y in the second equation with 2x + 3
2x + 3 = -x - 4
Collect like terms
3x = - 7
x = -7/3
If x = -7/3 then substitute that value in to the top equation to find y
y = 2(-7/3) + 3
y = -14/3 + 9/3
y = - 5/3